A BALL OF MASS m IS THROWN UPWARDS WITH A VELOCITY v . IF AIR EXERTS AN AVERAGE RESISTING FORCE r, THEN THE VELOCITY WITH WHICH IT RETURNS TO THE THROWER?

Grade:12

1 Answers

Arun
6 years ago
Dear Ragini
 
½ * m * u^2 = (mg+ r) * h
½ * m * v^2= (mg – r) * h
hence
v^2 / u^2 = (mg – r) / (mg + r)
v = u \sqrt(mg – r) / (mg + r)

Regards
Arun (askIITians forum expert)